Circular No. 7924

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dp IN NGC 7678 A. Klotz, Guitalens, France, reports the discovery of an apparent su-

pernova (mag ∼ 15.1 ± 0.4) on CCD images (limiting mag ∼ 17.0) taken on June 18.09 and 19.02 UT with a 0.20-m reflector. SN 2002dp is located at α ∼ 23^{h}28^{m}30^{s}.4, δ ∼ +22^{o}25′44′′ (equinox 2000.0), which is 30′′ east and 22′′ north of the nucleus of NGC 7678. The new object is not present on Palomar Sky Survey images or on an image taken on 2001 Sept. 11.94. W. Li, University of California, reports that SN 2002dp appears at mag ∼ 15.0 on a KAIT image taken on June 19.5 UT, from which the following position end figures were measured: 30^{s}.12, 38′′.8 (equinox 2000.0), or 30′′.8 east and 21′′.7 north of the nucleus of NGC 7678. A KAIT image taken on 2001 Dec. 28.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002dq IN NGC 7051 D. Hutchings and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on June 18.5 (mag ∼ 17.0) and 19.4 UT (mag ∼ 17.1). SN 2002dq is located at α = 21^{h}19^{m}49^{s}.95, δ = −8^{o}46′39′′.8 (equinox 2000.0), which is 20′′.4 west and 19′′.0 north of the nucleus of NGC 7051. A KAIT image taken on June 9.5 showed nothing at this position (limiting mag ∼ 18.0).

1E 2259+586 V. M. Kaspi, McGill University and Massachusetts Institute of Tech-

nology; and F. P. Gavriil, McGill University, report the detection with the PCA aboard RXTE of short x-ray bursts from the direction of the anomalous x-ray pulsar 1E 2259+586 (cf. IAUC 3745). The Chandra-determined position for this pulsar is α = 23^{h}01^{m}08^{s}.295, δ = +58^{o}52′44′′.45 (equinox 2000.0; uncertainty 0′′.60 in radius, 99-percent confidence; Hulleman et al. 2001, Ap.J. 563, L49). The bursting behavior was detected during a routine 15000-s duration monitoring observation that began on June 18.652 UT. In this observation, ∼ 45 bursts were detected, most of which had duration < 1–2 s, reminiscent of those seen in soft γ-ray repeaters. The mean 1-s flux per burst is ≈ 10^{−8} erg cm^{−2} s^{−1} (2–20 keV). Observations at other wavebands are encouraged.

2002 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green